OpenStudy (anonymous):

Please help! Solve the system by elimination: 2x^2+5y^2=98 2x^2-y^2=2

OpenStudy (anonymous):

we can do this together

OpenStudy (anonymous):

first subtract second equation from first equation

OpenStudy (anonymous):

6y^2=96

OpenStudy (anonymous):

i think u r not intrested :(

OpenStudy (anonymous):

No I'm back! Sorry... I had to log out and log back in. So what's the next step in solving using elimination? I really appreciate your help.

OpenStudy (anonymous):

2x^2-2x^2+5y^2+y^2=98-2 6y^2=96 y^2=96/6 y=root16 y=4

OpenStudy (anonymous):

now substitute y value in a equation to get x value

OpenStudy (anonymous):

x=3?

OpenStudy (anonymous):

yea :D

OpenStudy (anonymous):

So (4,3)? Thank you so much mathematicalpro :)

OpenStudy (anonymous):

yup :) always welcome :D

OpenStudy (anonymous):

sorry, its not (4,3) you should write x value first so,its (3,4)